OpenStudy (anonymous):

Can someone please help me find the x-intercepts and vertex of f(x)=x^2-7 and sketch a graph?

OpenStudy (kagıtucak):



OpenStudy (kagıtucak):

for intercepts: it is clear for y it is -7 for x take y=0 then x^2=7 x=\[\sqrt{7}\] or x=\[-\sqrt{7}\]

OpenStudy (anonymous):

i only need to find the x-intercepts, so the x-intercepts would actually just be √7 and -√7??

OpenStudy (kagıtucak):

yes exactly

OpenStudy (anonymous):

so I wouldnt factor the x^2 and 7 so that it is x(x-7) x=0 x=7?

OpenStudy (kagıtucak):

just look carefully x(x-7) is x^2-7x bu its equation is x^2-7 not -7x

OpenStudy (anonymous):

oh! so the vertex is (0,-7)?

OpenStudy (anonymous):